OpenStudy (anonymous):

find the laplace transform of f prime of t in terms of L(s) and use this result to find the laplace transform of cos(ax)

OpenStudy (anonymous):

i know L(s) of sin(ax), just don't know how to piece this equation together

OpenStudy (ybarrap):

$$ \large \frak{L}_s[f'(t)](s) =sF(s)\\ \frak{L}_s[\frac{d}{dt}\left(\frac{sin(at)}{a}\right)](s)\\ =s\frak{L}_s[\frac{sin(at)}{a}](s)\\ =\frac{s}{a}\frac{a}{a^2+s^2}\\ =\frak{L}_s^{-1}[\frac{s}{a^2+s^2}](t)\\ =\cos(at) $$